OpenStudy (anonymous):

HELP ASAP !!!!! To determine whether an element is actually gold, scientists will often determine the density of the element and compare it to the known density of gold (19.3 grams/cubic centimeter, or 19.3 g/cc). Samples less dense might be mixtures of many elements (not pure gold). A sample is found to have a mass of 500 g. The sample is a perfect 2 cm cube. Is this sample gold?

OpenStudy (debbieg):

How many cubic centimeters of volume are in a 2cm cube?

OpenStudy (debbieg):



OpenStudy (anonymous):

6?

OpenStudy (debbieg):

Hmmmm... do you know the formula for volume of a cube?

OpenStudy (anonymous):

no i don't.

OpenStudy (debbieg):

In general, volume of a rectangular solid is v=l*w*h In a cube, it is simply that l=w=h, so you get \(\large v=s^3\) where s is a side length (remember, the sides are all the same length).

OpenStudy (debbieg):

So here, you have s=2. So what is volume?

OpenStudy (anonymous):

6?

OpenStudy (debbieg):

What is s?

OpenStudy (anonymous):

speed...

OpenStudy (debbieg):

The formula is \(\large v=s^3\) or, if you prefer, you can just use \(\large v=s\cdot s \cdot s\)

OpenStudy (debbieg):

Noooo, I told you above... s is a side length of the cube.

OpenStudy (debbieg):

I'll repeat: In general, volume of a rectangular solid is v=l*w*h In a cube, it is simply that l=w=h, so you get \(\large v=s^3\) where s is a side length (remember, the sides are all the same length). so HERE, what is s=?

OpenStudy (anonymous):

side length

OpenStudy (debbieg):

Yes, that's what s is representing... I'm asking you, what is s=? What NUMBER is s=___???

OpenStudy (anonymous):

2

OpenStudy (debbieg):

Right. Now you need volume, and you know that \(\large v=s^3\) So \(\large v=2^3\) So what is the volume of the cube?

OpenStudy (anonymous):

8?

OpenStudy (debbieg):

Once you have volume (which will be in cubic centimeters) you are looking for g/cc, and you were already told grams in the problem. So you just need to divide: \[\Large \dfrac{ 500~grams }{ volume~cc }=g/cc\]

OpenStudy (debbieg):

Right, v=8. Now just take the 500 grams, divide by the 8 cc, and you'll get the g/cc.